OpenStudy (anonymous):

For i = , if 3i (2 + 5i) = x + 6i, then x = ?

OpenStudy (anonymous):

3i (2 + 5i) = x + 6i Foild \[6i+15i^2=x+6i\]\[15i^2+6i-6i-x=0\]\[15i^2-x=0\]\[i^2 =-1\]\[15(-1)-x=0\]\[-15-x=0\]\[x=-15\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with one more problem?